Please see the problem below, please be sure to read the numbers and question carefully and be sure to check for all unit conversions and double check your answer. Thank you!

Below the question is the outline of how to work the question out!

1) A charge, g1, of -8 micro-coulombs is located at the origin, a charge, q2, of -5 micro-coulombs is located at x- 0 cm, y +8.6 cm, a charge, q3, of +8 micro-coulombs is located at x = 0 cm, y =-2.6 cm, and a charge, q4, of +2 micro-coulombs is located at x-7.4 cm, y-0 cm, what is the magnitude of the total electric force on d2 (charge on the +y axis) in Newtons?


4, t) Charge 2 and 1: ks.l

五. Charge 2 and charge 3: ha 63 3 (t)

Charge 2 and charge 4: So, + ttx
Answer #1

F3 F4 Cosao

AD = distance between charge q3 and q2 = 8.6 + 2.6 = 11.2 cm = 0.112 m

AB = distance between charge q1 and q2 = 8.6 cm = 0.086 m

BC = distance between charge q4 and q2 = 7.4 cm = 0.074 m

Using Pythagorean theorem in triangle ABC

AC = sqrt(AB2 + BC2) = sqrt((0.086)2 + (0.074)2) = 0.1135 m

heta = tan-1(BC/AB) = tan-1(0.074/0.086) = 40.7 deg

F1 = force by charge q1 on q2 = k q1 q2 /AB2 = (9 x 109) (8 x 10-6) (5 x 10-6)/(0.086)2 = 48.7 N

F3 = force by charge q3 on q2 = k q3 q2 /AD2 = (9 x 109) (8 x 10-6) (5 x 10-6)/(0.112)2 = 28.7 N

F4 = force by charge q4 on q2 = k q4 q2 /AC2 = (9 x 109) (2 x 10-6) (5 x 10-6)/(0.1135)2 = 6.98 N

Net force along Y-direction is given as

Fy = F1 - F3 - F4 Cosheta = 48.7 - 28.7 - (6.98) Cos40.7 = 14.71 N

Net force along X-direction is given as

Fx = F4 Sinheta = (6.98) Sin40.7 = 4.55 N

magnitude of net force is given using Pythagorean theorem as

F = sqrt(Fx2 + Fy2) = sqrt((4.55)2 + (14.71)2) = 15.4 N